Shift report 20020109 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Jan 09 2002 - 07:58:53 EST

  • Next message: Flemming Videbaek: "regarding runs taken 1/8-1/9"

    ------------------------------------------------------------
    Date: Wed Jan  9 07:58:53 2002, Shift: 00:00-08:00
    Supervisor: ykl, with: 
    Log book#12 pages 30 - 33
    
    Shift Summary:
    ------------------------------------------------------------
    1. Continuous beam most of the night. Quality was poor
    for BRAHMS. Actual vertex determined by BB is shifted by
    70 cm from the nominal vertex. MCR worked on this problem
    for 2 hours by changing angles in both X and Y planes
    but finally gave up. They claim specialist has to come in.
    
    2. 
    
    Run 6384   MRS @40 deg, 4B;  Trig 3 has 96K
               FS  @8 deg, 1/8B; Trig 6 has 78K
    
    Run 6385   MRS @40 deg, 4B;  Trig 3 has 45K
               FS  @8 deg, 1/4B; Trig 6 has 40K
    
    3.  no new detector problems.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Jan 09 2002 - 07:59:32 EST